# Model Reconstruction Using Counterfactual Explanations: Mitigating the Decision Boundary Shift

### v1

@@ 3.2 An attack based on the proximity of counterfactuals to the target decision boundary
4 Experiments
We carry out a number of experiments to study the performance of our proposed strategy CCA. We include some results here and provide further details in Appendix B.
All the classifiers are neural networks unless specified otherwise and their decision boundaries are not necessarily convex. We compare the performance of our proposed attack CCA with the existing attack presented in (Aïvodji, Bolot, and Gambs 2020) that we refer to as “Baseline” for the case of one-sided counterfactuals. As the initial counterfactual generating method, we use an implementation of the Minimum Cost Counterfactuals (MCCF) by Wachter, Mittelstadt, and Russell (2017).
Results: A summary of the main experiments is provided below. Additional experiments are detailed in Appendix B.
(i) Visualizing the attack using synthetic data: First, the effect of the proposed loss function in mitigating the decision boundary shift is observed over a 2-D synthetic dataset. Fig. 6 presents the results. In the figure, it is clearly visible that the Baseline model is affected by a decision boundary shift. In contrast, the CCA model’s decision boundary closely approximates the target decision boundary. See Appendix B.2 for more details.
(ii) Comparing performance over four real-world dataset: We use four publicly available real-world datasets namely, Adult Income, COMPAS, DCCC, and HELOC (see Appendix B.1) for our experiments. Table 1 provides some of the results over four real-world datasets. We refer to Appendix B.2 (specifically Fig. 11) for additional results. In all cases, we observe that CCA has either better or similar fidelity as compared to Baseline.
@@ 4 Experiments
(iii) Studying effects of Lipschitz constants: We study the connection between the target model’s Lipschitz constant and the attack performance. Target model’s Lipschitz constant is controlled by changing the $L_{2}-$regularization coefficient, while keeping the surrogate models fixed. Results are presented in Fig. 7. Target models with a smaller Lipschitz constant are easier to extract. More details are provided in Appendix B.2.
@@
(iv) Studying different model architectures: We also consider different surrogate model architectures spanning models that are more complex than the target model to much simpler ones. Results show that when sufficiently close to the target model in complexity, the surrogate architecture plays a little role on the performance. Otherwise, the performance degrades. See Fig. 15 and Appendix B.2 for details. Furthermore, two situations are considered where the target model is not a neural network (Fig. 19 and Appendix B.2). In both scenarios, CCA surpasses the baseline.
@@
(v) Studying other counterfactual generating methods: Effects of counterfactuals being sparse, actionable, and realistic are observed. Sparse counterfactuals are generated by using $L_{1}-$norm as the cost function. Actionable counterfactuals are generated using DiCE (Mothilal, Sharma, and Tan 2020) by defining a set of immutable features. Realistic counterfactuals are generated by retrieving the 1-Nearest-Neighbor for a given query from the dataset. The results are given in Table 5 with additional details in Appendix B.2.
@@
(vi) Studying alternate loss functions: We explore using binary cross-entropy loss function directly with labels 0, 1 and 0.5, in place of the proposed loss. However, experiments indicate that this scheme performs poorly when compared with the CCA loss (see Fig. 20 and Appendix B.2).
@@
(vii) Validating Theorem 3.2: Empirical verification of the theorem is done through synthetic experiments, where the model has a spherical decision boundary since they are known to be more difficult for polytope approximation (Arya, Da Fonseca, and Mount 2012). Fig. 8 presents a log-log plot comparing the theoretical and empirical query complexities for several dimensionality values $d$. The empirical approximation error decays faster than $n^{-2/(d-1)}$ as predicted by the theorem (see Appendix B.3).
@@ 5 Conclusion
Our work provides novel insights that bridge explainability and privacy. We introduce novel model extraction strategies aided by counterfactuals, along with the analyses of the corresponding query complexities. The two attacks exploit (i) properties of closest counterfactuals; and (ii) the proximity of the counterfactuals to the target model’s decision boundary, respectively. Our work addresses an important knowledge gap in the existing literature by providing theoretical guarantees using random polytope theory. Furthermore, our attacks address the issue of decision boundary shift amidst a system providing one-sided counterfactuals. Experiments demonstrate a significant improvement in fidelity compared to the baseline method proposed in Aïvodji, Bolot, and Gambs (2020) for the case of one-sided counterfactuals. Results show that the attack performs well across different model types and counterfactual generating methods. Our findings also highlight an interesting connection between Lipschitz constant and vulnerability to model extraction.
Limitations and future work: Utilizing techniques in active learning in conjunction with counterfactuals is another problem of interest. Extending the results of this work for multi-class classification scenarios can also be explored. The relationship between Lipschitz constant and vulnerability to model extraction may have implications for future work on generalization, adversarial robustness, etc.
One limitation of our theoretical results is the assumption of a convex decision boundary. While this assumption may be satisfied in some settings (Amos, Xu, and Kolter 2017), it is not always the case. In addition, the derivations require the counterfactual generating mechanisms to provide the closest counterfactuals. Any relaxation to these assumptions are paths for exploration. We note that the attack proposed in Section 3.2 remains valid as long as the majority of counterfactuals lie closer to the decision boundary; they need not be the closest ones to the corresponding original instances.
6 Broader Impact
We demonstrate that one-sided counterfactuals can be used for perfecting model extraction attacks, exposing a potential vulnerability in MLaaS platforms. Given the importance of counterfactuals in explaining model predictions, we hope our work will inspire countermeasures and defense strategies, paving the way toward secure and trustworthy machine learning systems.
Appendix A Proofs
@@
We first have a look at Böröczky Jr and Reitzner (2004, Theorem 1 (restated as Theorem A.3 below)) from the polytope theory. Let ${\mathbb{M}}$ be a compact convex set with a second-order differentiable boundary denoted by $\partial{\mathbb{M}}$. Let $\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{n}$ be $n$ randomly chosen points on $\partial{\mathbb{M}}$, distributed independently and identically according to a given density $d_{\partial{\mathbb{M}}}$. Denote by $H_{+}(\boldsymbol{a}_{i})$ the supporting hyperplane of $\partial{\mathbb{M}}$ at $\boldsymbol{a}_{i}$. Assume $C$ to be a large enough hypercube which contains ${\mathbb{M}}$ in its interior.
@@
Theorem A.3 (Random Polytope Approximation, (Böröczky Jr and Reitzner 2004)).
@@
$\leq\underbrace{|\tilde{m}(\boldsymbol{x})-\tilde{m}(\boldsymbol{w})|}_{\leq\gamma_{1}||\boldsymbol{x}-\boldsymbol{w}||_{2}}+\underbrace{|m(\boldsymbol{x})-m(\boldsymbol{w})|}_{\leq\gamma_{2}||\boldsymbol{x}-\boldsymbol{w}||_{2}}$ (18)
$\leq(\gamma_{1}+\gamma_{2})||\boldsymbol{x}-\boldsymbol{w}||_{2}$ (19)
where the first inequality is a result of applying the triangle inequality and the second follows from the definition of Lipschitz continuity (Definition 2.7). ∎
@@ B.1 Details of Real-World Datasets
  • Adult Income: The dataset is a 1994 census database with information such as educational level, marital status, age and annual income of individuals (Becker and Kohavi 1996). The target is to predict “income”, which indicates whether the annual income of a given person exceeds \$50000 or not (i.e., $y=\mathds{1}[\text{income}\geq 0.5]$). It contains 32561 instances in total (the training set), comprising of 24720 from $y=0$ and 7841 from $y=1$. To make the dataset class-wise balanced we randomly sample 7841 instances from class $y=0$, giving a total effective size of 15682 instances. Each instance has 6 numerical features and 8 categorical features. During pre-processing, categorical features are encoded as integers. All the features are then normalized to the range $[0,1]$.
  • Home Equity Line of Credit (HELOC): This dataset contains information about customers who have requested a credit line as a percentage of home equity (FICO 2018). It contains 10459 instances with 23 numerical features each. Prediction target is “is_at_risk” which indicates whether a given customer would pay the loan in the future. Dataset is slightly unbalanced with class sizes of 5000 and 5459 for $y=0$ and $y=1$, respectively. Instead of using all 23 features, we use the following subset of 10 for our experiments; “estimate_of_risk”, “net_fraction_of_revolving_burden”, “percentage_of_legal_trades”, “months_since_last_inquiry_not_recent”, “months_since_last_trade”, “percentage_trades_with_balance”, “number_of_satisfactory_trades”, “average_duration_of_resolution”, “nr_total_trades”, “nr_banks_with_high_ratio”. ALl the features are normalized to lie in the range $[0,1]$.
  • Correctional Offender Management Profiling for Alternative Sanctions (COMPAS): This dataset has been used for investigating racial biases in a commercial algorithm used for evaluating reoffending risks of criminal defendants (Angwin et al. 2016). It includes 6172 instances and 20 numerical features. The target variable is “is_recid”. Class-wise counts are 3182 and 2990 for $y=0$ and $y=1$, respectively. All the features are normalized to the interval $[0,1]$ during pre-processing.
  • Default of Credit Card Clients (DCCC): The dataset includes information about credit card clients in Taiwan (Yeh 2016). The target is to predict whether a client will default on the credit or not, indicated by “default.payment.next.month”. The dataset contains 30000 instances with 24 attributes each. Class-wise counts are 23364 from $y=0$ and 6636 from $y=1$. To alleviate the imbalance, we randomly select 6636 instances from $y=0$ class, instead of using all the instances. All the attributes are numerical, and normalized to $[0,1]$ during pre-processing.
@@
Based on the experiments of Aïvodji, Bolot, and Gambs (2020) and Wang, Qian, and Miao (2022), we select $T,N=20$ and $S=100$, unless specified otherwise.
Visualizing the attack using synthetic data
This experiment is conducted on a synthetic dataset which consists of $1000$ samples generated using the make_moons function from the sklearn package. Features are normalized to the range $[0,1]$ before feeding to the classifier. The target model has 4 hidden layers with the architecture (10, 20, 20, 10). The surrogate model is 3-layered with the architecture (10, 20, 20). Each model is trained for 100 epochs. Since the intention of this experiment is to demonstrate the functionality of the modified loss function given in (3.2), a large query of size 200 is used, instead of performing multiple small queries. Fig. 6 shows how the original model extraction attack proposed by (Aïvodji, Bolot, and Gambs 2020) suffers from the boundary shift issue, while the model with the proposed loss function overcomes this problem. Fig. 10 illustrates the instances misclassified by the two surrogate models.

### v2

@@ 4 Experiments
(iii) Studying effects of Lipschitz constants: We study the connection between the target model’s Lipschitz constant and the CCA performance. Target model’s Lipschitz constant is controlled by changing the $L_{2}-$regularization coefficient, while keeping the surrogate models fixed. Results are presented in Fig. 15. Target models with a smaller Lipschitz constant are easier to extract. More details are provided in Appendix B.2.4.
@@
(iv) Studying different model architectures: We also consider different surrogate model architectures spanning models that are more complex than the target model to much simpler ones. Results show that when sufficiently close to the target model in complexity, the surrogate architecture plays a little role on the performance. See Appendix B.2.5 for details. Furthermore, two situations are considered where the target model is not a neural network in Fig. 16 and Appendix B.2.8. In both scenarios, CCA surpasses the baseline.
@@
(v) Studying other counterfactual generating methods: Effects of counterfactuals being sparse, actionable, realistic, and robust are observed. Sparse counterfactuals are generated by using $L_{1}-$norm as the cost function. Actionable counterfactuals are generated using DiCE (Mothilal et al., 2020) by defining a set of immutable features. Realistic counterfactuals (that lie on the data manifold) are generated by retrieving the 1-Nearest-Neighbor from the accepted side for a given query, as well as using the autoencoder-based method C-CHVAE (Pawelczyk et al., 2020). Additionally, we generate robust counterfactuals using ROAR (Upadhyay et al., 2021). We evaluate the attack performance on the HELOC dataset (Table 2). Moreover we observe the distribution of the counterfactuals generated using each method w.r.t. the target model’s decision boundary using histograms (Fig. 8). Additional details are given in Appendix B.2.6.
Table 2: Fidelity achieved with different counterfactual generating methods on HELOC dataset. Target model has hidden layers with neurons (20, 30, 10). Surrogate model architecture is (10, 20).
Fidelity over ${\mathbb{D}}_{\text{test}}$ Fidelity over ${\mathbb{D}}_{\text{uni}}$
CF method n=100 n=200 n=100 n=200
Base. CCA Base. CCA Base. CCA Base. CCA
MCCF L2-norm 91 95 93 96 91 93 93 95
MCCF L1-norm 93 95 94 96 89 92 91 95
DiCE Actionable 93 94 95 95 90 91 93 94
1-Nearest-Neightbor 93 95 94 96 93 93 94 95
ROAR (Upadhyay et al., 2021) 91 92 93 95 87 85 92 92
C-CHVAE (Pawelczyk et al., 2020) 77 80 78 82 90 89 85 78
Figure 8: Histograms of the target model’s predictions on different types of input instances. Counterfactual generating methods except MCCF with $L_{2}$ norm often generate counterfactuals that are farther inside the favorable region, hence having a target model prediction much greater than 0.5. We count all the query results across all the target models in the ensembles used to compute the average fidelities corresponding to each counterfactual generating method.
(vi) Comparison with DualCFX: DualCFX proposed by Wang et al. (2022) is a strategy that utilizes the counterfactual of the counterfactuals to mitigate the decision boundary shift. We compare CCA with DualCFX in Table 6, Appendix B.2.7.
@@
(vii) Studying alternate loss functions: We explore using binary cross-entropy loss function directly with labels 0, 1 and 0.5, in place of the proposed loss. However, experiments indicate that this scheme performs poorly when compared with the CCA loss (see Fig. 18 and Appendix B.2.10).
@@
(viii) Validating Theorem 3.2: Empirical verification of the theorem is done through synthetic experiments, where the model has a spherical decision boundary since they are known to be more difficult for polytope approximation (Arya et al., 2012). Fig. 20 presents a log-log plot comparing the theoretical and empirical query complexities for several dimensionality values $d$. The empirical approximation error decays faster than $n^{-2/(d-1)}$ as predicted by the theorem (see Appendix B.3).
@@ 5 Conclusion
Our work provides novel insights that bridge explainability and privacy through a set of theoretical guarantees on model reconstruction using counterfactuals. We also propose a practical model reconstruction strategy based on the analysis. Experiments demonstrate a significant improvement in fidelity compared to the baseline method proposed in Aïvodji et al. (2020) for the case of one-sided counterfactuals. Results show that the attack performs well across different model types and counterfactual generating methods. Our findings also highlight an interesting connection between Lipschitz constant and vulnerability to model reconstruction.
Limitations and Future Work: Even though Theorem 3.6 provides important insights about the role of query size in model reconstruction, it lacks an exact characterization of $k(\epsilon)$ and $v_{i}(\epsilon)$. Moreover, local Lipschitz continuity might not be satisfied in some machine learning models such as decision trees. Any improvements along these lines can be avenues for future work. Utilizing techniques in active learning in conjunction with counterfactuals is another problem of interest. Extending the results of this work for multi-class classification scenarios can also be explored. The relationship between Lipschitz constant and vulnerability to model reconstruction may have implications for future work on generalization, robustness, etc. Another direction of future work is to design defenses for model reconstruction attacks, leveraging and building on strategies for robust counterfactuals (Dutta et al., 2022; Hamman et al., 2024; Upadhyay et al., 2021; Black et al., 2022; Jiang et al., 2023; Pawelczyk et al., 2020).
Broader Impact: We demonstrate that one-sided counterfactuals can be used for perfecting model reconstruction. While this can be beneficial in some cases, it also exposes a potential vulnerability in MLaaS platforms. Given the importance of counterfactuals in explaining model predictions, we hope our work will inspire countermeasures and defense strategies, paving the way toward secure and trustworthy machine learning systems.
Appendix A Proof of Theoretical Results
@@
We first have a look at Böröczky Jr and Reitzner [2004, Theorem 1 (restated as Theorem A.3 below)] from the polytope theory. Let ${\mathbb{M}}$ be a compact convex set with a second-order differentiable boundary denoted by $\partial{\mathbb{M}}$. Let $\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{n}$ be $n$ randomly chosen points on $\partial{\mathbb{M}}$, distributed independently and identically according to a given density $d_{\partial{\mathbb{M}}}$. Denote by $H_{+}(\boldsymbol{a}_{i})$ the supporting hyperplane of $\partial{\mathbb{M}}$ at $\boldsymbol{a}_{i}$. Assume $C$ to be a large enough hypercube which contains ${\mathbb{M}}$ in its interior.
@@
Theorem A.3 (Random Polytope Approximation, [Böröczky Jr and Reitzner, 2004]).
@@
Consider the $i^{\text{th}}$ piece $m_{i}(\boldsymbol{x})$ of the classifier defined over ${\mathbb{U}}_{i}$. A part of the decision boundary exists within ${\mathbb{U}}_{i}$ only if $\exists\boldsymbol{x}\in{\mathbb{U}}_{i}$ such that $m_{i}(\boldsymbol{x})=0.5$. When it is the case, at the decision boundary,
$m(\boldsymbol{x})=0.5$ (15)
$\iff$ $\frac{1}{1+e^{-\ell_{i}(\boldsymbol{x})}}=0.5$ (16)
$\iff$ $e^{-\ell_{i}(\boldsymbol{x})}=1$ (17)
$\iff$ $\ell_{i}(\boldsymbol{x})=0$ (18)
$\iff$ $\boldsymbol{a}_{i}^{T}\boldsymbol{x}+b_{i}=0$ (19)
which represents a hyperplane restricted to ${\mathbb{U}}_{i}$. Moreover, the continuity of the $\ell(\boldsymbol{x})$ demands the decision boundary to be continuous across the boundaries of ${\mathbb{U}}_{i}$’s. This fact can be proved as follows:
Note that within each region ${\mathbb{U}}_{i}$, exactly one of the following three conditions holds:
  • (a)
[…]
  • (b)
[…]
  • (c)
[…]
In case when (c) holds for some region ${\mathbb{U}}_{i}$, the decision boundary within ${\mathbb{U}}_{i}$ is affine and it extends from one point to another on the region boundary. Now let ${\mathbb{U}}_{s}$ and ${\mathbb{U}}_{t},s,t\in\{1,\dots,q\},s\neq t$ be two adjacent regions sharing a boundary. Assume that ${\mathbb{U}}_{s}$ contains a portion of the decision boundary, which intersects with a part of the shared boundary between ${\mathbb{U}}_{s}$ and ${\mathbb{U}}_{t}$ (note that ${\mathbb{U}}_{i}$’s are closed and hence they include their boundaries). Let $\boldsymbol{x}_{0}$ be a point in the intersection of the decision boundary within ${\mathbb{U}}_{s}$ and the shared region boundary. Now, continuity of $\ell(\boldsymbol{x})$ at $\boldsymbol{x}_{0}$ requires $\ell_{t}(\boldsymbol{x}_{0})=\ell_{s}(\boldsymbol{x}_{0})=0$. Hence, condition (c) holds for ${\mathbb{U}}_{t}$. Moreover, this holds for all the points in the said intersection. Therefore, if such a shared boundary exists between ${\mathbb{U}}_{s}$ and ${\mathbb{U}}_{t}$, then the decision boundary continues to ${\mathbb{U}}_{t}$. Applying the argument to all ${\mathbb{U}}_{s}-{\mathbb{U}}_{t}$ pairs show that each segment of the decision boundary either closes upon itself or ends at a boundary of the unit hypercube. Hence, when taken along with the boundaries of the unit hypercube, the decision boundary is a collection of polytopes. ∎
A.4 Proof of Theorem 3.6
See 3.6
Proof.
Note that
[…]
Let $\mathcal{M}_{i}$ denote the event “$i^{\text{th}}$decision boundary cell does not have a counterfactual”. At the end of $n$ queries,
${\mathbb{P}}[\mathcal{M}_{i}]$ $=\prod_{j=1}^{n}\underbrace{{\mathbb{P}}[j^{\text{th}}\text{queryfalling outside of }{\mathbb{G}}_{m,g_{m}}({\mathbb{H}}_{i})]}_{=1-v_{i}(\epsilon)\text{ for uniform queries}}$ (23)
$=(1-v_{i}(\epsilon))^{n}.$ (24)
Therefore,
${\mathbb{P}}[\text{Reconstruction}]$ $=1-\sum_{i=1}^{k(\epsilon)}(1-v_{i}(\epsilon))^{n}$ (25)
$\geq 1-k(\epsilon)(1-v^{*}(\epsilon))^{n}\quad\left(\because v_{i}(\epsilon)\geq v^{*}(\epsilon)=\min_{j}v_{j}(\epsilon)\right).$ (26)
A.5 Proof of Theorem 3.10 and Corollary 3.11
See 3.10 See 3.11
Proof.
Note that from Theorem 3.6, with probability $p\geq 1-k(\epsilon)(1-v^{*}(\epsilon))^{n}$ at least one counterfactual exists within each decision boundary cell. When this is the case, we have
@@
$\leq\underbrace{|\tilde{m}(\boldsymbol{x})-\tilde{m}(\boldsymbol{w})|}_{\leq\gamma_{\tilde{m}}||\boldsymbol{x}-\boldsymbol{w}||_{2}}+\underbrace{|m(\boldsymbol{x})-m(\boldsymbol{w})|}_{\leq\gamma_{m}||\boldsymbol{x}-\boldsymbol{w}||_{2}}$ (29)
$\leq(\gamma_{m}+\gamma_{\tilde{m}})||\boldsymbol{x}-\boldsymbol{w}||_{2}$ (30)
$\leq\sqrt{d}(\gamma_{m}+\gamma_{\tilde{m}})\epsilon$ (31)
where the first inequality is a result of applying the triangle inequality and the second follows from the definition of local Lipschitz continuity (Definition 3.9). The final inequality is due to the availability of a counterfactual within each decision boundary cell, which ensures $||\boldsymbol{x}-\boldsymbol{w}||_{2}\leq\sqrt{d}\epsilon$. Corollary 3.11 follows directly from the second inequality, since the $\delta-$cover of $\boldsymbol{w}$’s ensure $||\boldsymbol{x}-\boldsymbol{w}||_{2}\leq\delta$
@@ Appendix B Experimental Details and Additional Results
All the experiments were carried-out on two computers, one with a NVIDIA RTX A4500 GPU and the other with a NVIDIA RTX 3050 Mobile.
@@ B.1 Details of Real-World Datasets
  • Adult Income: The dataset is a 1994 census database with information such as educational level, marital status, age and annual income of individuals [Becker and Kohavi, 1996]. The target is to predict “income”, which indicates whether the annual income of a given person exceeds \$50000 or not (i.e., $y=\mathds{1}[\text{income}\geq 0.5]$). It contains 32561 instances in total (the training set), comprising of 24720 from $y=0$ and 7841 from $y=1$. To make the dataset class-wise balanced we randomly sample 7841 instances from class $y=0$, giving a total effective size of 15682 instances. Each instance has 6 numerical features and 8 categorical features. During pre-processing, categorical features are encoded as integers. All the features are then normalized to the range $[0,1]$.
  • Home Equity Line of Credit (HELOC): This dataset contains information about customers who have requested a credit line as a percentage of home equity FICO [2018]. It contains 10459 instances with 23 numerical features each. Prediction target is “is_at_risk” which indicates whether a given customer would pay the loan in the future. Dataset is slightly unbalanced with class sizes of 5000 and 5459 for $y=0$ and $y=1$, respectively. Instead of using all 23 features, we use the following subset of 10 for our experiments; “estimate_of_risk”, “net_fraction_of_revolving_burden”, “percentage_of_legal_trades”, “months_since_last_inquiry_not_recent”, “months_since_last_trade”, “percentage_trades_with_balance”, “number_of_satisfactory_trades”, “average_duration_of_resolution”, “nr_total_trades”, “nr_banks_with_high_ratio”. All the features are normalized to lie in the range $[0,1]$.
  • Correctional Offender Management Profiling for Alternative Sanctions (COMPAS): This dataset has been used for investigating racial biases in a commercial algorithm used for evaluating reoffending risks of criminal defendants [Angwin et al., 2016]. It includes 6172 instances and 20 numerical features. The target variable is “is_recid”. Class-wise counts are 3182 and 2990 for $y=0$ and $y=1$, respectively. All the features are normalized to the interval $[0,1]$ during pre-processing.
  • Default of Credit Card Clients (DCCC): The dataset includes information about credit card clients in Taiwan Yeh [2016]. The target is to predict whether a client will default on the credit or not, indicated by “default.payment.next.month”. The dataset contains 30000 instances with 24 attributes each. Class-wise counts are 23364 from $y=0$ and 6636 from $y=1$. To alleviate the imbalance, we randomly select 6636 instances from $y=0$ class, instead of using all the instances. Dataset has 3 categorical attributes, which we encode into integer values. All the attributes are normalized to $[0,1]$ during pre-processing.
@@
Based on the experiments of Aïvodji et al. [2020] and Wang et al. [2022], we select $T=20,50,100$; $N=20,8,4$ and $S=100,50$, in different experiments. We note that the exact numerical results are often variable due to the multiple random factors affecting the outcome such as the test-train-attack split, target and surrogate model initialization, and the randomness incorporated in the counterfactual generating methods. Nevertheless, the advantage of CCA over the baseline attack is observed across different realizations.
B.2.1 Visualizing the attack using synthetic data
This experiment is conducted on a synthetic dataset which consists of $1000$ samples generated using the make_moons function from the sklearn package. Features are normalized to the range $[0,1]$ before feeding to the classifier. The target model has 4 hidden layers with the architecture (10, 20, 20, 10). The surrogate model is 3-layered with the architecture (10, 20, 20). Each model is trained for 100 epochs. Since the intention of this experiment is to demonstrate the functionality of the modified loss function given in (3), a large query of size 200 is used, instead of performing multiple small queries. Fig. 7 shows how the original model reconstruction proposed by Aïvodji et al. [2020] suffers from the boundary shift issue, while the model with the proposed loss function overcomes this problem. Fig. 10 illustrates the instances misclassified by the two surrogate models.